OpenStudy (anonymous):

need help with... 2u times 2u cubed times (vcubed) squared

OpenStudy (anonymous):

Does that mean \(2u \times 2u^3 \times (v^3)^2\)?

OpenStudy (anonymous):

yea, and so what's the question?

OpenStudy (anonymous):

(4u^4)(v^6)

OpenStudy (anonymous):

when multiplying variables like u^3 and u^1 you add the 1 and the 3 to get 4 (V^3)(v^3)=V^3 squared add 3 and 3 to get 6

OpenStudy (anonymous):

@Dalvoron yes it does & thanks guys ! i really apprieciate all the help!

OpenStudy (anonymous):

The general rules for combining powers in this case are \[x^a \times x^b = x^{a+b}\]\[(x^a)^b=x^{a \times b}\]